# Large Language Models in Qualitative Research: Can We Do the Data Justice?

### v1

@@ 6. Limitations
While we recruited participants across disciplines, our participant group does not cover all perspectives on qualitative research and its values. Disciplines that are even more situated than those represented in our participant group, such as ethnography, may be at greater risk of diminishment by LLM involvement, and those perspectives may not be represented by the voices highlighted here. We believe the more contextually situated, slower, and interpretivist qualitative epistemologies are, the more LLMs are likely to be at odds with or threaten their research values. We also recruited participants regardless of their level of engagement with LLMs as users or as researchers. This elicited important findings about current use by non-experts, important intuitions from qualitative researchers regardless of their level of use, and insights from experts on responsible AI who avoid LLMs. However, some of these researchers, by virtue of not being experts on LLMs and their potential flaws, may have presented a different set of concerns than experts, particularly on the nature of risks from LLMs or their potential capabilities under different conditions of use. The fact our participants were mostly PhD students also likely shapes the set of concerns elicited, level of engagement with these tools, and interest in future use of LLMs in research. We also note our participants largely did not surface experience or opinions on simulated research subjects or automated interviewing, which are garnering significant attention in other research circles.
7. Implications
In this section, we consider the practical implications of our findings for two audiences: LLM-curious qualitative researchers, who are wondering if and how qualitative data work can be conducted ethically and effectively, and future qualitative data analysis tool builders, who are contending with how to build tools that facilitate these experiences.
7.1. Guidelines for researchers using LLMs in qualitative work
Based on our interview findings and prior work, we outline a set of guidelines for qualitative researchers interested in using LLMs in their research process. Qualitative researchers are currently already using and likely expanding their use of LLMs in research. The goal of these guidelines is to scaffold an approach for conscientious use of these tools beyond the current moment for researchers who may be considering whether and how to use them.
7.1.1. Intentionally select tools
Most participants who had used LLMs referred to ChatGPT, which is the most popular LLM interface currently available. However, ChatGPT is not an interface designed to be used for qualitative research. Understanding what data a tool was trained on, whether it responds well to research-related prompts, and whether it handles data appropriately. While there are not currently bespoke LLM models for qualitative research, they might exist in the future, and dedicated analysis platforms are incorporating LLM-based features.
7.1.2. Follow venue and institutional policies
Several venues have created policies for appropriate use and disclosure of LLMs in research, including the ACM and *ACL conferences (noa, [n. d.]; Chairs, [n. d.]). For example, the ACM allows use of generative AI, but a manuscript cannot be “primarily the result of the tool’s generative capabilities,” and authors must disclose generative AI use in the paper. The publishing venue Sage has similar guidelines, allowing use for editing but banning reviewers from inappropriately using generative AI (Sage, 2024). Follow your community’s guidelines, and consider how you might shape them going forward.
Beyond adhering to institutional guidelines mentioned above, when using LLMs as part of an analysis method or for analysis, disclose use fully and be explicit about where models were used, including which model was used and any fine-tuning, optimization, or prompt engineering which was performed. You may wish to engage in additional disclosure, particularly for uses in qualitative research that may have shaped original contributions forming theory on the data, like the generation of themes from the data using an LLM. On the other hand, in the absence of explicit venue or university guidelines, you may decide that light editing is appropriate without disclosure, such as the use of Grammarly to lightly edit texts, if tools did not shape the theoretical contribution or framing of the work.
7.1.3. Validate outputs for your use case
Prioritize validation of LLM outputs in proportion to the significance of a result. Our results raise concerns that the core intellectual contribution of qualitative research may be subject to the temptation to trust an LLM “collaborator.” Regardless of methodology, you must remain responsible for your process and results. You should choose methods of validating the use of LLMs according to your use case and methodological values, selecting tools and methods that are transparent and disclosing your process as much as possible. When using LLMs to produce small-$N$ outputs, such as a single interview protocol or translation, you should ensure these are carefully checked and reviewed. If using LLMs to produce large-$N$ outputs, such as applying qualitative codes at scale or deploying personalized interview guides, you should have a plan to validate a subsample of the data and check performance on your task. Furthermore, you should consider whether more transparent methods than using LLMs for your chosen task exist, and if so, prioritize using them, since current lack of transparency and explainability of LLM outputs makes transparency in the research process more difficult. Whether or not LLM outputs can be validated according to an existing standard may depend on the chosen epistemology. For example, depending on your research values, you may decide not to use LLM-generated codes, if you decide that your research places value on human-generated perspectives on coding instead. You should also consider whether you may be committing accidental acts of academic plagiarism by relying on LLM-generated outputs (Bhardwaz and Kumar, 2023) across the research pipeline, and should take responsibility for checking outputs for any copied content accordingly.
7.1.4. Protect participant data
Qualitative researchers must protect participant data and anonymity. Across the board, but especially if using a closed source or proprietary model, take extreme care not to share personally identifiable information or compromise the data of at-risk populations. Consider whether hosting open source models locally, privately, and securely is feasible based on your project constraints as well as suitable for your use case. If you plan to use a proprietary model, carefully consider how you might reliably anonymize data and whether participants have been given the chance to provide explicit consent to have their data shared with third-party models for analysis. Avoid LLM interfaces which will or could use your participants’ data as training data, given risks to their interests, without their consent. Consider providing explicit justification if using a proprietary model for scientific purposes, as recommended by  Palmer et al. (2024). If using an external software tool, such as NVivo or Atlas.ti, that incorporates LLM-based functions, check whether the tool is using a proprietary model that may change your assessment of its appropriate use with your data, either according to university policy or IRB standards. As a matter of process, update your consent forms for participants to accurately reflect any intended use of LLMs in the research and implications for participants. This should include details on how participant privacy will be protected (through using anonymized data on local models, for example, in contrast to companies that may use participant data to train future models).
7.1.5. Consider a model’s training data when evaluating potential harms
The major LLMs available today, such as ChatGPT, Claude, or Gemini, have been trained on large swaths of data on the Internet (Kandpal et al., 2023), but commercial models largely lack transparency on their data sources (Dubey et al., 2024; Soldaini et al., 2024). They perform best when applied to data that is closer to their training data, as our participants noted, but may appear confident and — on the surface — plausible, despite a lack of transparency into their mechanisms. However, the details of qualitative research processes are typically not published on the Internet to preserve participant privacy, nor is there a large amount of qualitative research data in general, so it is unknown but unlikely that major models have been trained in ways that are well-fitted to qualitative research praxis or to specialized groups of participants for tasks like coding data. They also may perform especially poorly on data concerning non-Western participants, non-English speaking participants, or people from otherwise non-dominant backgrounds. If your research, like much qualitative research, focuses on these groups, using LLMs could misalign with the goals of your research.
7.1.6. Consider potential risks to your sensemaking process
When considering how to use LLMs, consider whether the tool is doing the data and your research questions justice. For example, consider whether you may be using an LLM as a way to spend less time with the data, whether you are introducing a limited or biased view on the data, and whether you could become overreliant on the tool. All researchers should consider whether using a tool that changes the sensemaking process diminishes or enhances it, particularly depending on one’s research values.
We have identified the urgent need for transparent tools that center qualitative research use cases. Participants already use chat-based proprietary LLM interfaces like ChatGPT for qualitative research tasks, while dedicated analysis tools, like NVivo, do not currently serve comparable roles or have comparable features. Conversely, general-purpose chat-based LLM interfaces are not designed as qualitative research tools. As such, we outline some use cases that present opportunity areas for the development of tools for qualitative researchers that leverage LLMs.
Based on our findings, tool builders ought to emphasize 1) interactivity, 2) data exploration, and 3) qualitative coding in future qualitative research tools. Participants consistently mentioned the useful nature of the interactivity LLMs afford through a chat-based interface, a new modality compared to standard analysis software tools. Participants highlighted that chat interactions help them ideate, obtain feedback, and that they would appreciate making requests for data retrieval or formatting in natural language as is increasingly possible  (Izacard and Grave, 2021; Gao et al., 2023b). There is a gap in tools that facilitate these use cases seamlessly integrated with other desired uses. Second, researchers also perceive an opportunity to use LLMs to explore their data or see it in a new way. LLMs could help researchers see their data through a multiplicity of lenses, “support serendipity” (Jiang et al., 2021) for qualitative researchers and ideally widen insight rather than homogenize it. Lastly, qualitative researchers are interested in using LLMs to conscientiously assist in the coding process. Tools could also help scaffold the qualitative coding process for beginners, but some worry that tools may diminish ”essential friction” (P15) as one learns coding. Tools should help researchers check if the LLM is implementing their desired definitions and codes correctly. Across all opportunity areas, implementing LLM-based features also present risks which must be navigated carefully, such as the risks of anthropomorphizing AI through interactivity, becoming overreliant on it, decreasing researcher agency, reducing novelty, or the risk of broader misalignment with qualitative research values.
Researchers want to use tools that align with their values, and some implementations could empower researchers, while others may risk dispossessing them or participants. Below are a few principles for tool designers when considering how and if to create tools that responsibly address these opportunity areas:
7.2.1. Design for participant privacy
Proprietary LLMs have sparked warranted concerns regarding participant privacy. Some researchers mention closed source models as a potential solution, but most are unable to implement solutions to this problem themselves. Making it easy to host open-source models for some qualitative research functions would help qualitative researchers feel more confident about using LLMs in qualitative research. Tools that incorporate LLMs should be explicit about when and how they call external APIs, how they handle privacy, how they expect users to deanonymize data before using tools, how they scrub or plan to delete data, and whether they use inputs for training models.
7.2.2. Design for validation
Different epistemologies will have different ideas about what it means to validate outputs. Part of designing for researchers to feel agency and trust is giving them ways to validate whether the use of LLMs in a process gives useful and valid outputs and make evaluation transparent to let researchers feel empowered to make choices and feel confident in their decisions throughout the research process. This includes designing features to understand or probe outcomes and suggestions, as well as designing for reproducibility of outcomes when possible. Features that require researchers to review outputs depending on the task may help accomplish this goal.
7.2.3. Design for agency
Researchers want to maintain agency over if, when, and how they are assisted or influenced by AI tools in their research process. One way to design for agency is to give researchers ways of opting in or out of AI-based features within platforms, facilitating researchers’ own development of theory from the data before providing assistance. Designing for agency also involves designing against overreliance on or anthropomorphism of these technologies.
7.2.4. Design for researcher context
Researchers often wish to bring in their own context when performing qualitative analysis, including their past experiences, the texts that influence their work, and the theories in which they ground their work. Tools for qualitative researchers should be designed assuming researchers have unique perspective to contribute to the data.
7.2.5. Design for deep engagement with data
Qualitative researchers worry that LLM-based tools could distance them from data rather than deepen their relationship with it, but tools can be explicitly designed to elicit deeper engagement with the data. Many qualitative researchers expressed interest in using LLMs to see their data through another lens, but found that insights and themes produced on their data are generic or repetitive. Tools can explore LLMs as a way to produce more unique ways of understanding their data, drawing researchers’ attention to less predictable features of the data rather than predictable ones, or helping researchers examine or challenge their existing theories with direct evidence from the data, ideally deepening researchers’ understanding of the data rather than diminishing it.
7.2.6. Design to consider participant perspectives and interests
Researchers are concerned about the variable performance LLMs have across contexts, knowledge domains, cultures, and languages. Tools should work to facilitate better alignment between LLMs and diverse subject populations. When possible, tool builders could give the option to use models that better reflect a target, like models that specialize in a particular language or knowledge domain, and facilitate investigating where there might be errors or bias, in alignment with the recommendation to design for agency and validation.
@@ 8. Conclusion
Involving LLMs in qualitative research presents potential opportunities for augmenting research with new capabilities, but also surfaces tensions that show some types of LLM involvement may misalign with the values of qualitative research. Careful engagement with these problems as researchers, and thoughtful tool design from developers, can help navigate these tensions in order to both align with research values and support researchers in their intended use case. ChatGPT spurred new ideas for researchers regarding how their research processes might be augmented, but it should not serve as either a model or an endpoint for what is possible for LLMs in research tooling. Instead, we can integrate concerns, opportunities, and lessons to keep qualitative research values at the center of our work. The guidelines and design principles we suggest can help qualitative researchers and those in related fields, including journalism and many social sciences, center human perspectives to keep doing participants and their data justice as we navigate a new era of tooling.
9. Acknowledgements
We thank James Eschrich for their formative contributions to the scope, literature, and analysis in this piece. We also thank Mor Naaman, Emily Tseng, and Louise Barkhuus for their perspectives on this work. We thank the participants of the “LLMs as Research Tools” workshop at CHI 2024 for their contributions that informed the direction of this work.

### v2

@@ 7. Limitations
While we recruited participants across disciplines, our participant group does not cover all perspectives on qualitative research and its values. Disciplines that are even more situated than those represented in our participant group, such as ethnography, may be at greater risk of diminishment by LLM involvement, and those perspectives may not be represented by the voices highlighted here. The fact that our participants were mostly PhD students also likely shapes the findings, since younger people have more positive views of AI (Stein et al., 2024), and senior researchers are more likely to have ethical concerns about AI research use (Liao et al., 2024). While our participants represent a diverse group of people studying various populations, there was a heavy bias for American universities, which may miss key perspectives from non-Western institutions. Despite these limitations, our study highlights an important population wrestling with incorporating transformative technologies into their research processes for the first time.
@@ 8. Conclusion
LLMs are poised to continue proliferating and permeating research. While LLMs present potential opportunities for assistance in some research tasks, they also surface potential tensions with the values of qualitative research. In this climate, we document a sociotechnical breakdown: the functionality of tooling is insufficient to reliably preserve participant privacy, the guidelines and norms lag behind, and the social context increasingly incentivizes the broad use of AI across surfaces. Active, intentional consideration from researchers on issues like privacy and model bias in advance of task selection, coupled with thoughtful tool design, can help to maintain agency as researchers decide appropriate uses for LLMs. We synthesized current uses, tensions, and ethical challenges into a set of choices for the intentional incorporation of LLMs in qualitative research. We hope that this paper will empower qualitative researchers to leverage LLMs confidently, and even creatively, for their work if they choose.
Acknowledgements.
We thank James Eschrich for their formative contributions to the scope, literature, and analysis in this piece. We also thank Mor Naaman, Emily Tseng, Louise Barkhuus, and students and staff at the MIT Center for Constructive Communication for their perspectives on this work. We thank the participants of the “LLMs as Research Tools” workshop at CHI 2024 for their contributions that informed the direction of this work. We acknowledge that we used ChatGPT and Elicit to find specific works of related literature, which we then verified from source documents.